# Yufei Bo, Shuo Shao, Meixia Tao

### v1

@@ IV-A Experiment Settings
IV-A1 Dataset
Our experiments are conducted on the CIFAR100 dataset[40], which includes 50000 training images and 10000 test images. The resolution of each image is $32\times 32$. All images are classified into 20 super-categories, and images in each super-category are further classified to 5 sub-categories[40]. The 20-category classification label of each image stands for its coarse-grained semantic source $\mathbf{S}_{1}$ while the 100-category classification label of each image stands for its fine-grained semantic source $\mathbf{S}_{2}$.
@@ IV-A2 NN Architecture and Hyperparameters
The transmitter design has been presented in Section III. For decoders, we adopt Spinal-net[41] for coarse classification as well as fine classification. The semantic decoders for image recovery at both receivers have the same NN architecture, which consists of Resnet blocks combined with the depth-to-space operation to perform the upsampling. Table II presents the details of the NN architecture we use in this paper. We employ the Adam optimizer for the training. The detailed learning rate (LR) schedule is shown in Table III. The experiments are conducted using two super-constellations: 4QAM×$\times$4QAM and 4QAM×$\times$16QAM. We set the value of the power allocation factor $a$ such that the super-constellations form rectangular QAM constellation. Specifically, for 4QAM×$\times$4QAM, we set $a=0.80$, and for 4QAM×$\times$16QAM, we set $a=0.76$. All the experiments are performed on Intel Xeon Silver 4214R CPU, and 24 GB Nvidia GeForce RTX 3090 Ti graphics card with Pytorch powered with CUDA 11.4.
(a) Image recovery vs. $r$.
(b) Classification accuracy vs. $r$.
Figure 4: Performances of the DeepSCM scheme and the benchmarks at varying transmission rates, with the super-constellation set as 4QAM×$\times$16QAM.
@@ IV-A3 Benchmarks
We compare the proposed DeepSCM scheme with the conventional unstructured coded modulation (CM) scheme using rectangular $M$-QAM modulation, where only one semantic encoder is utilized at the transmitter. Additionally, we compare our DeepSCM scheme with classical separation-based source and channel coding scheme for a single receiver, where JPEG2000 and ideal capacity-achieving channel code are employed (abbreviated as “JPEG2000+Capacity”). In the CM scheme, the NN architecture of the semantic encoder, modulator and the two receivers are the same as their counterparts in our proposed DeepSCM scheme. We employ three different training methods for the CM scheme as three benchmarks.
  • CM Joint Training: In this scheme, the transmitter and the two receivers are jointly trained. Specifically, the loss function is set to be a weighted sum of the distortion measures of both receivers, which means that the transmitter balances the feedback from both receivers.
  • CM Trained with Rx1: Following the training method in [34], in this scheme, the transmitter is jointly trained with Receiver 1 in the absence of Receiver 2. After the transmitter is trained and fixed, the decoders at Receiver 2 is then trained to achieve its best possible performance. Note that this scheme serves as the performance upperbound for Receiver 1 under the CM scheme.
  • CM Trained with Rx2: This training scheme follows the same idea of “CM Trained with Rx1” but exchanges the role of Receiver 1 and Receiver 2. Similarly, it serves as the performance upperbound for Receiver 2 under the CM scheme.
(a) Image recovery vs. $r$.
(b) Classification accuracy vs. $r$.
Figure 5: Performances of the DeepSCM scheme and the benchmarks at varying transmission rates, with the super-constellation set as 4QAM×$\times$4QAM.
Figure 6: Visual examples of the image recovered by different schemes for Receiver 1 and Receiver 2 with $M=64$ and $r=\frac{1}{6}$.
@@ IV-A4 Performance Metrics
For the coarse and fine classification tasks, we use classification accuracy to evaluate the performance. For the image recovery task, we use peak-signal-to-noise ratio (PSNR) to denote the performance, which is defined as
$\text{PSNR}=10\log\frac{\text{MAX}^{2}}{\text{MSE}},$ (19)
where MAX is the maximum possible pixel value of the image. We define the transmission rate $r$ as the ratio between the number of channel uses and the dimension of the images, i.e.,
$r=\frac{n}{k},$ (20)
where for CIFAR100, $k=32\times 32\times 3$.
@@ IV-B Performances at Varying Transmission Rates
In this subsection, we compare the performances of the DeepSCM scheme and the benchmarks when the transmission rate varies from $\frac{1}{24}$ (128 channel uses) to $\frac{1}{4}$ (768 channel uses). If not specified otherwise, we set the channel SNR of Receiver 1 as $-5$ dB and that of Receiver 2 as 20 dB.
Fig. 4 focuses on the case when a 4QAM×$\times$16QAM super-constellation is used for our proposed DeepSCM scheme. Fig. 4(a) illustrates the PSNR performance of image recovery of Receiver 1 and Receiver 2 versus different transmission rates. Overall we can observe that the PSNRs of the proposed DeepSCM scheme for both receivers are very close to their respective upperbound. Specifically, for Receiver 1, the performance of the proposed scheme nearly coincides with the performance upperbound, the CM trained with Rx1 scheme, outperforming the CM trained with Rx2 scheme by more than 1 dB. Notice that we omit the performance of JPEG2000$+$Capacity for Receiver 1, since this scheme completely breaks down under the poor channel condition of Receiver 1. For Receiver 2, the performance of the proposed scheme approaches the performance upperbound, the CM trained with Rx2 scheme, and coincides with it when $r$ is high. The proposed scheme outperforms the CM joint training scheme by 0 dB to 1 dB, and outperforms the CM trained with Rx2 scheme by more than 2 dB. Moreover, the proposed scheme has a great advantage over the conventional separate coding scheme, particularly at lower transmission rates.
@@
(a) The image recovery vs. SNR.
(b) Classification accuracy vs. SNR.
Figure 7: The performance comparison of the DeepSCM scheme and the CM joint training scheme at different channel SNR values of Receiver 2.
(a) The convergence curve of image recovery.
(b) The convergence curve of classification accuracy.
Figure 8: The convergence curve of the DeepSCM scheme with different $a$.
(a) Image recovery vs. $a$.
(b) Classification accuracy vs. $a$.
Figure 9: Performance of the DeepSCM scheme with varying $a$.
@@
Figure 10: The empirical distributions of the 4QAM×$\times$16QAM super-constellation points with different PAF values.
Fig. 6 displays visual examples of the image recovered by different schemes for Receiver 1 and Receiver 2. We can observe that for Receiver 2, the proposed DeepSCM scheme, along with the CM trained with Rx2 scheme, attains the highest quality in recovered images. The images recovered by the CM joint training scheme and the CM trained with Rx1 scheme appear more blurred, indicating their inferior performances. Moreover, compared with the images recovered by the DeepSCM scheme, those recovered from JPEG 2000 compression display a loss of sharpness around high-contrast edges and exhibit a blocky structure. These JPEG artifacts negatively affect the visual quality of the images. In contrast to the images of Receiver 2, those of Receiver 1 exhibit lower quality due to its poor channel condition. Furthermore, since Receiver 1 has a smaller channel capacity, it cannot decode the images encoded by the JPEG 2000 compression algorithm and the capacity-achieving channel coding of Receiver 2, which shows the limitations of the conventional separate coding scheme in a broadcast channel.
All in all, simulation results show that the proposed superposition code structure is indeed efficient in degraded broadcast channels. It alleviates the conflict between diverse semantic information granularity requirements of the two receivers with different channel conditions, improving the performance of both receivers instead of sacrificing the performance of one receiver to improve that of the other.
IV-C Performances versus Channel Disparity
In this subsection, we compare the performance of our proposed DeepSCM scheme and that of the CM joint training scheme at varying levels of channel disparity, namely different SNR gaps between Receiver 1 and Receiver 2. Specifically, we fix the channel SNR of Receiver 1 as $-5$ dB, and vary the channel SNR of Receiver 2 from 0 dB to 20 dB. The transmission rate is set as $\frac{1}{6}$ (512 channel uses).
Fig. 7 shows the performances of the DeepSCM scheme and the benchmarks at different levels of channel disparity. We tune the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ so that the two schemes have the same performance in the classification task, as shown in Fig. 7(b), and we compare their performances in the image recovery task, illustrated in Fig. 7(a). Two important observations can be made. First, our scheme has an increasing performance advantage over the joint training scheme as the channel disparity increases. For example, when using a 4QAM×$\times$4QAM super-constellation, the performance of our scheme at Receiver 2 and that of the joint training scheme almost coincide when the channel disparity is lower than 15 dB. When the channel disparity rises up to 25 dB, our scheme outperforms the joint training scheme by 0.3 dB. Additionally, we can see that higher modulation order will further enlarge the performance gap between our scheme and the joint training scheme. For example, when the channel disparity is 25 dB, or equivalently, the SNR of Receiver 2 is 20 dB, the performance advantage of our scheme can reach 0.6 dB when a 4QAM×$\times$16QAM super-constellation is used, 0.3 dB higher than the performance advantage when a 4QAM×$\times$4QAM super-constellation is used.
In conclusion, the stimulation results show that our scheme exhibits a performance advantage over the CM joint training scheme that correlates positively with the communication capability gap between the two receivers. Specifically, when there is a larger disparity in channel SNR between the receivers or when the system supports a higher modulation order, our scheme yields greater benefits compared to the joint training scheme due to its superposition coding structure.
IV-D Impact of PAF
@@
Since power allocation between different receivers is an important issue in classical superposition code, in this subsection, we investigate the impact of the hyperparameter PAF by varying $a$ from 0.5 to 0.9. The experiment is conducted using the 4QAM×$\times$16QAM super-constellation. The channel SNR for Receiver 1 is set to -5 dB, while Receiver 2 has an SNR of 20 dB. We fix the parameter $r$ at $\frac{1}{6}$, equivalent to 512 channel uses. Intuitively, there should be a performance trade-off between the two receivers. Allocating more power to the inner constellation points will improve the performance of Receiver 1, and vice versa.
@@
Fig. 8 shows the convergence curve of the DeepSCM scheme with different values of $a$, illustrating their impact on each training stage of the DeepSCM scheme. Fig. 8(a) shows the convergence curve of image recovery. We can indeed observe a performance trade-off between the two receivers during the first two stages. In Stage 1 when Receiver 1 is trained, larger values of $a$ result in better performance of Receiver 1, which is attributed to the fact that the superposed outer constellation points serve as additional noise. In Stage 2 when Receiver 2 is trained, larger values of $a$ (e.g., $a=0.9$) lead to worse performance for Receiver 2. However, this intuitive performance trade-off disappears during the fine-tuning stage, where larger values of $a$ also lead to decreased performance for Receiver 1. This indicates that Receiver 1 can still decode the outer constellation to some extent due to the powerful adaptability of NNs to channel noise. The influence of $a$ on the performance of classification in Fig. 8(b) is less obvious, yet similar observations can be made.
@@
Fig. 9 further shows the performance of the DeepSCM scheme with varying values of $a$. Fig. 9(a) shows the performance of image recovery of the two receivers. The performance of Receiver 1 decreases with increasing $a$. For Receiver 2, not only too large an $a$ causes its performance degradation, but too small an $a$ also leads to worse performance. This is because with smaller $a$, the outer constellation points will overlay, making it harder for Receiver 2 to decode them. In Fig. 9(b) we can also observe that extreme PAF values cause a decline in performance. Given our hierarchical-source scenario, it is important to note that the results differ from the scenario of broadcasting independent messages to two receivers. In the latter case, increasing $a$ enhances the performance of one receiver while diminishing that of the other.
@@
Overall, these observations offer principles for determining the PAF: $a$ should be chosen to prevent the overlay of outer constellation points, and the interval between outer constellation points should not be too small. Therefore, it is best to choose moderate PAF values to achieve the overall best performance. Demonstrating that moderate PAF values are most reasonable significantly simplifies the selection of PAF. We can simply set $a$ such that the super-constellation forms rectangular QAM constellation, as in Section IV-B and Section IV-C.
@@
In Figure 10, we illustrate the empirical distributions of the 4QAM×$\times$16QAM super-constellation points with varying PAF values. It can be noted that the outer constellation points overlap with smaller $a$. Moreover, it is evident that due to the poor channel condition of Receiver 1, constellation points with lower power are less likely to be generated.

### v2

@@ IV-A Experiment Settings
IV-A1 Dataset and Channel
Our experiments are conducted on the CIFAR100 dataset[41], which includes 50000 training images and 10000 test images. The resolution of each image is $32\times 32$. All images are classified into 20 super-categories, and the images in each super-category are further classified to 5 sub-categories, resulting a total of 100 categories [41]. The 20-category classification label of each image stands for its coarse-grained semantic information $\mathbf{S}_{1}$ while the 100-category classification label of each image stands for its fine-grained semantic information $\mathbf{S}_{2}$. Throughout the simulation, we consider two-user Gaussian broadcast channels. If not specified otherwise, the SNR of receiver 1 is fixed at -5 dB while the SNR of receiver 2 varies between 0 dB and 20 dB. Receiver 1 intends to recover the original image as well as to perform the 20-category coarse classification while Receiver 2 intends to recover the original image and to perform the 100-category fine classification.
@@ IV-A2 NN Architecture and Hyperparameters
The transmitter design has been presented in Section III. The semantic decoders for image recovery at both receivers have the same NN architecture, which consists of Resnet blocks combined with the depth-to-space operation to perform the upsampling. Spinal-net[42] is adopted for both coarse and fine classification. Table II presents the details of the NN architecture we use in this paper. We employ the Adam optimizer for the training. The learning rate (LR) schedule is shown in Table III. The experiments are conducted using two super-constellations: 4QAM×$\times$4QAM and 4QAM×$\times$16QAM. We set the value of the PAF $\alpha$ such that the super-constellations form rectangular QAM constellations. Specifically, for 4QAM×$\times$4QAM, we set $\alpha=0.80$, and for 4QAM×$\times$16QAM, we set $\alpha=0.76$. All the experiments are performed on Intel Xeon Silver 4214R CPU, and 24 GB Nvidia GeForce RTX 3090 Ti graphics card with Pytorch powered with CUDA 11.4.
@@ IV-A3 Benchmarks
We compare the proposed DeepSCM with the conventional unstructured coded modulation (CM) scheme, where only one semantic encoder with rectangular M-QAM modulation is utilized at the transmitter. In the CM scheme, the NN architectures of the semantic encoder, modulator, and the two receivers are the same as their counterparts in DeepSCM. Three different training methods shall be used for the CM scheme as three different benchmarks.
  • CM Trained Jointly: In this scheme, the transmitter and the two receivers are jointly trained, where the loss function is set to be a weighted sum of the distortion measures of both receivers. This training scheme tries to balance the performance of the two receivers.
  • CM Trained with Rx1 (Performance Upperbound for Rx1): Following the training method in [21], in this scheme, the transmitter and Receiver 1 are first jointly trained in the absence of Receiver 2. After the transmitter is trained and fixed, Receiver 2 is then trained alone to achieve its best possible performance. Note that this scheme serves as the performance upperbound for Receiver 1 under the CM scheme.
  • CM Trained with Rx2 (Performance Upperbound for Rx2): This training scheme follows the same approach as “CM trained with Rx1” but exchanges the role of Receiver 1 and Receiver 2. Similarly, it serves as the performance upperbound for Receiver 2 under the CM scheme.
Additionally, we compare our DeepSCM scheme with classical separate source and channel coding (SSCC) schemes. Three SSCC schemes are considered as baselines, namely: JPEG2000 with capacity-achieving channel code (abbreviated as “JPEG2000+Capacity”), Better Portable Graphics (BPG) with capacity-achieving channel code (“BPG+Capacity”), and a more practical scheme, BPG with low density parity check (LDPC) codes (“BPG+LDPC”). The LDPC codes are from the widely-used IEEE 802.11n standard, with block length 648 bits for rate $\frac{1}{2}$, $\frac{2}{3}$ and $\frac{3}{4}$. As a two-user broadcast channel is considered in this paper, the SSCC baselines are implemented with two multiple-access strategies. One is the single-user (SU) strategy where the coding is designed for one user only, and the other is time division multiple access (TDMA), where each user occupies half of the channel uses and the coding is specifically designed for that user during its allocated channel uses.
(a) Performance comparison for Rx1: (left) classification accuracy; (center) PSNR; (right) MS-SSIM.
(b) Performance comparison for Rx2: (left) classification accuracy; (center) PSNR; (right) MS-SSIM.
Figure 4: Performance of different schemes at varying transmission rates, with 4QAM×$\times$16QAM super-constellation.
@@ IV-A4 Performance Metrics
For the coarse and fine classification tasks, we use classification accuracy to evaluate the performance. For the image recovery task, we utilize the pixel-wise performance metric PSNR and the perceptual metric MS-SSIM[43] to evaluate the performance.
We define the transmission rate $r$ as the ratio between the number of channel uses and the dimension of the images, i.e.,
$r=\frac{n}{d},$ (22)
where for CIFAR100 one has $d=32\times 32\times 3$.
(a) Performance comparison for Rx1: (left) classification accuracy; (center) PSNR; (right) SSIM.
(b) Performance comparison for Rx2: (left) classification accuracy; (center) PSNR; (right) SSIM.
Figure 5: Performance of different schemes at varying transmission rates, with 4QAM×$\times$4QAM super-constellation.
Figure 6: Visual examples of the image recovered by different schemes for Rx1 and Rx2 with $M=64$ and $r=0.167$.
(a) Performance comparison for Rx1: (left) PSNR; (right) classification accuracy.
(b) Performance comparison for Rx2: (left) PSNR; (right) classification accuracy.
Figure 7: Performance comparison at different channel SNRs of Rx2. The SNR of Rx1 is constantly set as -5 dB.
@@ IV-B Performances at Varying Transmission Rates
In this subsection, we compare the performance of DeepSCM and the benchmarks when the transmission rate varies from 0.04 (128 channel uses) to 0.25 (768 channel uses). If not specified otherwise, we set the channel SNR of Receiver 1 as $-5$ dB and that of Receiver 2 as 20 dB.
@@
Fig. 5 focuses on the cases where a 4QAM×$\times$4QAM super-constellation is used. Similar to the observations from Fig. 4, our proposed scheme can achieve the upperbound performance of both Receiver 1 and 2 simultaneously on all three measures, and achieves a better balance between the performances of the two receivers than the “BPG/JPEG2000+capacity” baselines.
Finally, Fig. 6 displays visual examples of the image recovered by different schemes for Receiver 1 and Receiver 2. Note that since BPG’s performance is generally better than JPEG2000’s, we omit the latter in this comparison. It can be observed that the proposed scheme achieves relatively strong recovery performance for both receivers, while other schemes sacrifice the performance of one receiver for the benefit of the other.
IV-C Performances versus Channel Disparity
In this subsection, we compare the performance of DeepSCM with the “CM trained jointly” scheme and the “BPG+LDPC” scheme using TDMA at varying channel disparity levels, namely different SNR gaps between Receiver 1 and Receiver 2. Specifically, we fix the channel SNR of Receiver 1 as $-5$ dB, and vary the channel SNR of Receiver 2 from 0 dB to 20 dB. The transmission rate is set as 0.167 (512 channel uses).
Fig. 7(a) and Fig. 7(b) show the performance comparison for Receiver 1 and Receiver 2, respectively, at different channel disparity levels. As a side note, we tune the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ so that DeepSCM and “CM trained jointly” have the same performance in the classification task, and we compare their performances in the image recovery task. Three important observations can be made. First, comparing the proposed scheme and the “BPG+LDPC” scheme, we can see that for Receiver 1, DeepSCM outperforms “BPG+LDPC” by a significant margin; and for Receiver 2, DeepSCM outperforms the “BPG+LDPC” scheme in image recovery performance in low SNR regions ($<11$ dB) and constantly outperforms it in classification accuracy across all evaluated SNRs. Moreover, the proposed scheme avoids the cliff effect in the “BPG+LDPC” scheme, which is due to the failure of the channel code in adverse channel conditions with SNRs lower than the code’s target SNR value. Second, our scheme has an increasing PSNR performance advantage over “CM trained jointly” as the channel disparity increases. For example, when using a 4QAM×$\times$4QAM super-constellation, the performance of our scheme for Receiver 2 and that of the “CM trained jointly” scheme almost coincide when the channel disparity is lower than 15 dB. When the channel disparity rises up to 25 dB, our scheme outperforms the “CM trained jointly” scheme by 0.25 dB. Third, we can see that higher modulation order will further enlarge the performance gap between our scheme and the “CM trained jointly” scheme. For example, when the channel disparity is 25 dB (the SNR of Receiver 2 is 20 dB), the PSNR performance gain of our scheme can reach 0.7 dB when a 4QAM×$\times$16QAM super-constellation is used, which is 0.45 dB higher than the gain obtained when a 4QAM×$\times$4QAM super-constellation is used.
Figure 8: Performance of Receiver 1 trained at particular SNRs while tested across varying channel SNRs: (left) PSNR; (right) classification accuracy. Channel use is set as 512 with 4QAM×$\times$16QAM. The channel SNR of Receiver 2 is set as 10 dB for both training and testing.
Figure 9: Performance of Receiver 2 trained at particular SNRs while tested across varying channel SNRs: (left) PSNR; (right) classification accuracy. Channel use is set as 512 with 4QAM×$\times$16QAM. The channel SNR of Receiver 1 is set as -5 dB for both training and testing.
(a) Convergence curve of image recovery.
(b) Convergence curve of classification accuracy.
Figure 10: Convergence of the DeepSCM scheme with varying PAF values.
(a) Image recovery vs. $\alpha$.
(b) Classification accuracy vs. $\alpha$.
Figure 11: Performance of the DeepSCM scheme with varying PAF values.
@@ IV-C Performances versus Channel Disparity
Figure 12: Empirical distributions of the 4QAM×$\times$16QAM super-constellation points with varying PAF values.
In conclusion, the simulation results demonstrate that our scheme outperforms the “CM trained jointly” scheme, with its performance advantage increasing as the communication capability gap between the two receivers widens. Moreover, DeepSCM avoids the cliff effect experienced by the “BPG+LDPC” scheme and outperforms it in low SNR regions.
IV-D Robustness to Variations in Channel Conditions
This subsection investigates the robustness of the proposed scheme to variations in channel conditions. To this end, the NNs in the system are trained at particular SNRs and tested over varying SNRs. We denote the training SNR as $\text{SNR}_{\text{train}}$ and testing SNR as $\text{SNR}_{\text{test}}$.
Fig. 8 and Fig. 9 illustrate the performance of Receiver 1 and Receiver 2 respectively at varying $\text{SNR}_{\text{test}}$s. For image recovery at both receivers, there is a graceful PSNR degradation when $\text{SNR}_{\text{test}}<\text{SNR}_{\text{train}}$, and performance improves but gradually saturates when $\text{SNR}_{\text{test}}$ keeps increasing. For classification, the accuracy performance is much more robust to the variation of $\text{SNR}_{\text{test}}$ around $\text{SNR}_{\text{train}}$, compared to the image recovery.
Additionally, the system attains its peak performance at a specific SNR when optimized for it, and a bigger gap between $\text{SNR}_{\text{train}}$ and $\text{SNR}_{\text{test}}$ results in worse performance at this $\text{SNR}_{\text{test}}$. This observation suggests a deployment strategy for the DeepSCM scheme in dynamic wireless conditions, i.e., storing multiple pretrained models trained with different channel SNRs at the transmitter and the receivers and then deploying the one with the training SNR closest to the current channel condition.
IV-E Impact of PAF
@@
Power allocation between different receivers is an important issue in classical superposition code. In this subsection, we investigate the impact of the hyperparameter PAF by varying the $\alpha$ in (4) from 0.5 to 0.9. The experiment is conducted using the 4QAM×$\times$16QAM super-constellation. The channel SNR for Receiver 1 is set to -5 dB, while Receiver 2 has an SNR of 20 dB. We fix the transmission rate $r$ at 0.167 (equivalent to 512 channel uses). Intuitively, there should be a performance trade-off between the two receivers. Allocating more power to the inner constellation points will improve the performance of Receiver 1, and vice versa.
@@
Fig. 10 shows the convergence of DeepSCM with different values of $\alpha$. From Fig. 10(a), we can indeed observe a trade-off in PSNR performance between the two receivers during the first two stages. In Stage 1, larger values of $\alpha$ result in better performance for Receiver 1, which is attributed to the fact that the outer constellation points serve as additional noise. In Stage 2, larger values of $\alpha$ (e.g., $\alpha=0.9$) lead to worse performance for Receiver 2. This intuitive performance trade-off, however, disappears during the fine-tuning stage, where larger values of $\alpha$ also lead to decreased performance for Receiver 1. This indicates that Receiver 1 can still decode the outer constellation to some extent due to the powerful adaptability of NNs to channel noises. The influence of $\alpha$ on classification accuracy in Fig. 10(b) is less obvious, yet similar observations can be made.
@@
Fig. 11 further compares the performance of the DeepSCM scheme with varying values of $\alpha$. From the PSNR performance comparison in Fig. 11(a), it is shown that the performance of Receiver 1 decreases with increasing $\alpha$. This is because, during the fine-tuning stage, it can better decode the outer constellation with a smaller $\alpha$. For Receiver 2, not only too large an $\alpha$ causes its performance degradation, but too small an $\alpha$ also leads to worse performance. This is because with a smaller $\alpha$, the outer constellation points will overlay, making it harder for Receiver 2 to decode them. The impact of PAF values on classification performance in Fig. 11(b) is observed to be less significant. Finally, it’s worth noting that these results differ from broadcasting independent messages to two receivers, given our hierarchical-source scenario. In the former case, increasing $\alpha$ improves the performance of one receiver while reducing that of the other.
@@
Overall, these observations offer guidelines for determining the PAF: $\alpha$ should be chosen to prevent the overlay of outer constellation points, and the interval between outer constellation points should not be too small. Therefore, it is best to choose moderate PAF values to achieve the overall best performance. The observation that moderate PAF values are most reasonable significantly simplifies the selection of PAF. We can simply set $\alpha$ such that the super-constellation forms rectangular QAM constellation, as in Section IV-B and Section IV-C.
@@
In Figure 12, we illustrate the empirical distributions of the 4QAM×$\times$16QAM super-constellations with varying PAF values. It can be noted that the outer constellation points overlap with a smaller $\alpha$. Moreover, it is evident that due to the poor channel condition of Receiver 1, constellation points with lower power are less likely to be generated.
IV-F Complexity Analysis
[…]
Table V compares the inference delay of different schemes. The proposed DeepSCM scheme and the CM scheme exhibit similar average encoding and decoding time, and they are both faster than BPG, while JPEG2000 shows the shortest inference delay. It is important to note that here the inference time of JPEG2000 and BPG refers to the source coding time only, and does not include the time for channel coding and task execution, i.e., classification in this work, which takes on average 7.2 ms.
TABLE IV: The number of FLOPs and model parameters with $n=512$ channel uses.
FLOPs (G) Params (M)
DeepSCM Tx 4.3 102.7
Rx1 1.5 24.7
Rx2 1.5 24.9
CM Tx 2.2 81.3
Rx1 1.5 24.7
Rx2 1.5 24.9
TABLE V: Inference delay comparison.
Encoding Time (ms) Decoding Time (ms)
SCM 16.9 11.0
CM 16.3 11.0
JPEG2000 3.8 0.6
BPG 28.3 18.5